Two balls are drawn. If there are 5 yellow balls and 8 violet balls, let Y represent yellow balls and let V represent violet balls. Let X be the random variable presents the number of violet balls. Find the probability of getting one violet ball.

1
Expert's answer
2022-03-25T06:15:26-0400

"P\left( X=1 \right) =\frac{8\cdot 5}{C_{13}^{2}}=0.512821"